# Toward Robust, Reproducible, and Widely Accessible Intracranial Language Brain-Computer Interfaces: A Comprehensive Review of Neural Mechanisms, Hardware, Algorithms, Evaluation, Clinical Pathways and Future Directions

### 03 Mar 2026 v1

@@ 7 Future Directions
On the data side, a foundational priority is the development of larger, multi-language, and clinically representative datasets to improve generalization across populations and linguistic contexts [135, 71]. Current evidence is dominated by English-speaking participants and highly controlled tasks, which limits external validity [116]. Expanding data collection to include tonal and morphologically diverse languages, broader age ranges, and heterogeneous etiologies of speech impairment will improve both scientific rigor and translational relevance [57].
@@ 8 Conclusion
This review summarizes the current trajectory of intracranial language BCIs for speech restoration across neural mechanisms, recording technologies, decoding algorithms, evaluation strategies, and translational design. Overall, the field is progressing from proof-of-concept demonstrations toward more clinically meaningful communication systems, but key bottlenecks remain. At the neural level, intracranial evidence supports a mixed organization with partially somatotopic articulatory coding, dorsal–ventral stream specialization, and frequency-specific dynamics across overt, mimed, and imagined speech. For covert/imagined speech, decoding is increasingly feasible in constrained settings, yet clinically reliable open-vocabulary decoding remains unresolved. At the systems level, recording modalities continue to present trade-offs in coverage, resolution, invasiveness, and long-term robustness. Multi-month stability has been demonstrated for both ECoG and intracortical approaches, but recalibration burden and setup dependence remain important barriers to routine longitudinal use. In parallel, advances in sequence models, transformers, and articulatory intermediates have improved decoding performance, while language priors (including LLM-based components) may enhance fluency and error correction; however, intent preservation, latency, bias, and privacy risks require continued safeguards and prospective monitoring.

### 16 Mar 2026 v2

@@ 7 Future Directions
On the data side, a foundational priority is the development of larger, multi-language, and clinically representative datasets to improve generalization across populations and linguistic contexts [137, 73]. Current evidence is dominated by English-speaking participants and highly controlled tasks, which limits external validity [118]. Expanding data collection to include tonal and morphologically diverse languages, broader age ranges, and heterogeneous etiologies of speech impairment will enhance both scientific rigor and translational relevance [59].
@@ 8 Conclusion
This review summarizes the current trajectory of intracranial language BCIs for speech restoration across neural mechanisms, recording technologies, decoding algorithms, evaluation strategies, and translational design. Overall, the field is progressing from proof-of-concept demonstrations toward more clinically meaningful communication systems, but key bottlenecks remain. At the neural level, intracranial evidence supports a mixed organization with partially somatotopic articulatory coding, dorsal–ventral stream specialization, and frequency-specific dynamics across overt, mimed, and imagined speech. For covert and imagined speech, decoding is increasingly feasible in constrained settings, yet clinically reliable open-vocabulary decoding remains unresolved. At the systems level, recording modalities continue to present trade-offs in coverage, resolution, invasiveness, and long-term robustness. Multi-month stability has been demonstrated for both ECoG and intracortical approaches, but recalibration burden and setup dependence remain important barriers to routine longitudinal use. In parallel, advances in sequence models, transformers, and articulatory intermediates have improved decoding performance, while language priors (including LLM-based components) may enhance fluency and error correction; however, intent preservation, latency, bias, and privacy risks require continued safeguards and prospective monitoring.